Subtract. Write your answer in simplest form.
(y-2/y+3) minus (y/2y+6)


Answer by checkley75(3666)

(y/2y)+6=1/2+6=6.5
y-2/y+3
-6.5
-------------------------
y-2/y-3.5
(y^2-2)/y-3.5